In the displacement method,a convex lens is placed in between an object and a screen. If the magnification in the two positions are m1 and m2 and the displacement of the lens between the two positions is x,then the focal length of the lens is
(1)x/m1+m2 (2)x/m1-m2 (3) x(m1+m2)² (4)x/(m1-m2)²
1
Expert's answer
2020-05-15T16:10:12-0400


For magnification "m_1":


"i_1=-m_1o_1,\\
\frac{1}{f}=\frac{1}{o_1}+\frac{1}{i_1}=\frac{1}{o_1}-\frac{1}{m_1o_1}=\frac{1}{o_1}(1-1/m_1)."

By analogy for magnification "m_2":


"\frac{1}{f}=\frac{1}{o_2}(1-1/m_2)"

From the figure we see that


"x=o_1-o_2\rightarrow o_2=o_1-x,"

substitute this:


"\frac{1}{f}=\frac{1}{o_1-x}(1-1/m_2),"

solve the system that consists of the last equation and this one:


"\frac{1}{f}=\frac{1}{o_1}(1-1/m_1)."

Simple math gives us the solution:


"f=\frac{x}{m_1-m_2}."